OpenStudy (anonymous):

Is 2(x-1)^2 = y+ 3 already in Standard Conic Form??

OpenStudy (anonymous):

i think this is in standard form, but i would not bet on it let me see if i can find the definition

OpenStudy (anonymous):

Wpouldnt it be (x-1)^2=y+3

OpenStudy (anonymous):

Ops i ment (x-1)^2=1/2(y+3)

OpenStudy (anonymous):

according to what i just found, standard form is \[a(x-h)^2=y-k\]

OpenStudy (anonymous):

so the coefficient goes in front of the square term

OpenStudy (anonymous):

soo (x-1)^2=1/2(y+3) ?

OpenStudy (anonymous):

no leave the 2 in front of the square term

OpenStudy (anonymous):

i found it in a quick google search \[a(x-h)^2=y-k\] it says, so leave the 2 in front of the \((x-1)^2\) term